OpenStudy (anonymous):

Henry is taking a medicine for a common cold. His body metabolizes the medicine over a certain time interval. The table below shows the amount of medicine f(t), in mg, that was present in Henry's body after time t:

OpenStudy (anonymous):

t (hours) 1 2 3 4 5 f(t) (mg) 282 265.08 249.18 234.22 220.17

OpenStudy (anonymous):

Greg was administered 200 mg of the same medicine. The amount of medicine in his body f(t) after time t is shown by the equation below: \[f(t)=200(0.88)^{t}\] Which statement best describes the rate at which Henry’s and Greg's bodies eliminated the medicine?

OpenStudy (anonymous):

Which statement best describes the rate at which Henry’s and Greg's bodies eliminated the medicine?

OpenStudy (anonymous):

(A) Henry's body eliminated the antibiotic faster than Greg's body. (B) Henry's body eliminated the antibiotic at the same rate as Greg's body. (C) Henry's body eliminated the antibiotic at half of the rate at which Greg's body eliminated the antibiotic. (D) Henry's body eliminated the antibiotic at one-fourth of the rate at which Greg's body eliminated the antibiotic.

OpenStudy (anonymous):

@TheSmartOne @hartnn @iambatman @paki I need help with this question plz

OpenStudy (paki):

@hartnn any idea about this question please...?

OpenStudy (anonymous):

@iGreen @sammixboo @mathmath333 @Michele_Laino I really need help on this question guys. Please help me out.

OpenStudy (michele_laino):

I'm trying...

OpenStudy (anonymous):

Thank you!

OpenStudy (michele_laino):

data referring to Greg, are: 1--->176 2--->154.88 3--->136.3 4--->120 5--->105.5

OpenStudy (anonymous):

okay I understand that

OpenStudy (michele_laino):

I used the formula above

OpenStudy (anonymous):

\[176=200(0.88)^{1}\]

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

And so on by changing the exponent until 5

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

I understand those parts but I just don't know what to do next

OpenStudy (michele_laino):

now, we can try to find the percentage reduction of the antibiotic in both Greg and Henry bodies. So in the case of Greg, we have to solve this equation \[176-\frac{ p }{ 100 }*176=105.5\]

OpenStudy (michele_laino):

p is the percentage rate of elimination of the antibiotic from Greg

OpenStudy (anonymous):

okay. what's next?

OpenStudy (michele_laino):

we have to solve that equation for p, namely we have to find the value of p

OpenStudy (anonymous):

how do we do that?

OpenStudy (michele_laino):

first I subtract from both sides 105.5, namely: \[176-\left( \frac{ p }{ 100 }*176 \right)-105.5=105.5-105.5\]

OpenStudy (michele_laino):

so, we have: \[70.5-\left( \frac{ p }{ 100 }*176 \right)=0\]

OpenStudy (michele_laino):

then: \[70.5=\frac{ p }{ 100 }*176\]

OpenStudy (michele_laino):

finally: \[p=\frac{ 70.5 }{ 176 }*100=40\]

OpenStudy (michele_laino):

so Greg body eliminates the antibiotic at a rate of about 40%

OpenStudy (anonymous):

okay I get that

OpenStudy (michele_laino):

Now, we have to solve the same equation for Henry, using the data referring to Henry. So we have to solve this equation: \[282-\frac{ q }{ 100 }*282=220.17\]

OpenStudy (michele_laino):

like before, I subtract from both sides 220.17: \[282-\left( \frac{ q }{ 100 }*282 \right)-220.17=220.17-220.17\]

OpenStudy (michele_laino):

so: \[61.83-\left( \frac{ q }{ 100 }*282 \right)=0\]

OpenStudy (michele_laino):

then: \[\frac{ q }{ 100 }*282=61.83\]

OpenStudy (michele_laino):

finally: \[q=\frac{ 61.83*100 }{ 282 }=22\]

OpenStudy (anonymous):

so its the third option?

OpenStudy (michele_laino):

so we can say that Henry body eliminates the antibiotic at a rate of about 22%, namely about the half of the rate of Greg

OpenStudy (anonymous):

bcuz 22/40 is equal to 0.55 which is closer to 0.5

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Thank YOU SO MUCHH! You were very helpful.

OpenStudy (michele_laino):

Thank you!